A saturated solution of magnesium fluoride , MgF2, was prepared by dissolving solid MgF2 in water. The concentration of Mg2+ ion in the solution was found to be 1.18
Answer #1

Ksp =(Mg+2)(2F-)^2

Ksp = (1.18*10^-3)(2*1.18*10^-3)^2

Ksp = 6.57*10^-9 ( mole/litre)^3
Answer #2

Ksp =(Mg+2)(2F-)^2

Ksp = (1.18*10^-3)(2*1.18*10^-3)^2

Ksp = 6.57*10^-9